Suppose that you buy a euro to US$ currency futures contract, where each relates to 1 million euros. You buy the contract at a price of $1.3468/€ (i.e. you commit to sell 1,000,000 × $1.3468 = $1,346,800 and receive 1 million euros). Over the next four days the contract closes at the following prices: $1.3465/€, $1.3443/€, $1.3434/€, and $1.3534/€. What would be your payments to, or withdrawals from, the margin account?

–$500, –$1,800, –$1,100, +$10,000

–$300, –$2,200, –$900, +$10,000

+$300, +$2,200, +$900, –$10,000

None of the answers are correct.
Answer #1

Payment or withdrawal = (Price on particular day - Previous price)*Number of Euros

Day 1 = (1.3465-1.3468)*1,000,000 = -$300

Day 2 = (1.3443-1.3465)*1,000,000 = $-$2,200

Day 3 = (1.3434 - 1.3443)*1,000,000 = -$900

Day 4 = (!.3534 - 1.3434)*1,000,000 = +$10,000

Hence, the answer is

–$300, –$2,200, –$900, +$10,000